# Full Line Code Completion: Bringing AI to Desktop

### v1

@@
Abstract.
In recent years, several industrial solutions for the problem of multi-token code completion have appeared, each making a great advance in the area but mostly focusing on cloud-based runtime and avoiding working on the end user’s device.
@@ Abstract.
Our online evaluation shows that the usage of the tool leads to 1.5 times more code in the IDE being produced by code completion. The described solution was initially started with a help of researchers and was bundled into two JetBrains’ IDEs—PyCharm Pro and DataSpell—at the end of 2023, so we believe that this work is useful for bridging academia and industry, providing researchers with the knowledge of what happens when complex research-based solutions are integrated into real products.
1. Introduction
@@
Code completion is among the most attractive features in integrated development environments (IDEs) (Murphy et al., 2006; Amann et al., 2016), since it helps programmers to avoid wasting time on manually typing all of the code characters and thus speeds up any kind of software engineering activity. Depending on the specific IDE, the complexity of code completion (and thus its quality) varies a lot: it can be a simple engine that just completes strings that a developer has already used in the current file, or it can use more complex approaches like static code analysis to predict possible tokens that a programmer may currently need. In any case, built-in completion engines usually only work in a single-token mode, i.e., only one identifier is generated (e.g., a variable or a method name).
@@
In the past years, several products for multi-token code completion have appeared on the market, including Copilot (Microsoft, 2024a), TabNine (TabNine, 2024), Codeium (Codeium, 2024), Amazon CodeWhisperer (Amazon, 2024), aiXcoder (aiXcoder, 2024), and others. Most of such tools are cloud-based, so they miss a wide niche of customers, which includes users beyond a firewall, as well as users with limited internet access. While TabNine and aiXcoder support fully local models and data querying, they lack integration features and do not implement code correctness checks. This still leaves an open market for code completion engines that are local-based and ensure a proper user experience in an IDE.
@@
In recent years, machine learning techniques were studied and applied for the different aspects of the completion problem (Chen et al., 2021; Bibaev et al., 2022; Wang et al., 2020; Svyatkovskiy et al., 2021), including the generation of the completion sequences itself. Since source code can be interpreted as sequential data, natural language processing (NLP) methods have found their application for the task (Hindle et al., 2016). Specifically, Transformer-based (Vaswani et al., 2017) neural networks significantly outperform other architectures in the task of sequential data generation (Radford et al., 2019; Brown et al., 2020; Raffel et al., 2020; Hellendoorn et al., 2019).
@@
At JetBrains, we built the Full Line Code Completion (FLCC) feature (JetBrains, 2024c) — a multi-token code completion engine that utilizes a Transformer-based neural network in its core and suggests only correct code. The feature is implemented in the form of a plugin for the IntelliJ Platform (JetBrains, 2024d; Kurbatova et al., 2021) (an open-source platform for building IDEs), and is bundled into two JetBrains IDEs—PyCharm Pro (JetBrains, 2024e) and DataSpell (JetBrains, 2024b)—where it is enabled for Python by default. The feature has been available since versions 2023.3 that were released in December 2023. It works fully locally, i.e. all computations, from code completion invocation to the showing of suggestion, happen on the end user’s device. Following modern industrial standards, the plugin is implemented in Kotlin, however, it utilizes an additional native server that is run locally and is implemented in C++. The whole model training pipeline is written mostly in Python, reusing the Kotlin pipeline only for the data processing step.
@@
In this paper, we describe our approach to building this Full Line Code Completion feature. Starting from the study of existing datasets, data collection, and the pre-processing with multiple tokenization specifics, we describe the details of the neural network training process and the end-user runtime, including a modified beam search algorithm, post-processing features, and user experience (UX) choices. While we cannot provide a full replication package of the described product feature, we share multiple conceptual insights we derived from our development experience. The focus of the paper is on the PyCharm Pro and DataSpell 2023.3 release state of the Full Line Code Completion, for which 100M parameters GPT-2 (Radford et al., 2019) model was used and inline completion (i.e., gray text completion) mode was implemented as a part of the IntelliJ Platform API (see Figure 1b). While preparing for the release, we conducted rigorous offline evaluation, as well as online evaluation in the form of A/B testing. Our experiments showed that the ratio of the completed code (a percentage of symbols of code written with code completion among all written code) increased 1.5 times for users with Full Line Code Completion compared to users with standard code completion only.
@@
  • We provide an in-depth exploration of the implementation of Full Line Code Completion for PyCharm Pro and DataSpell 2023.3 releases in December 2023.
@@
The rest of the paper is organized as follows. In Section 2, we describe the related work in the field. Then, in Section 3, we describe our approach in detail, highlighting various problems we encountered and how we solved them. Section 4 presents the results of our A/B testing experiments, and Section 5 discusses these results and our lessons learned. Finally, in Section 6, we describe the threats to the validity of our work, and in Section 7, we conclude the paper.
2. Background and related work
2.1. Background
2.1.1. Code completion
@@
Code completion, a prevalent feature across various IDEs (Murphy et al., 2006; Amann et al., 2016), is aimed at optimizing the work of a developer, not only by speeding up code typing but also by facilitating API exploration. Common in modern IDEs, single-token, or standard, code completion suggests individual tokens via a lookup window (see Figure 1a). This feature, relying on static code analysis, generates suggestions from multiple candidate providers, which include both built-in IDE modules and external plugins.
@@
In addition to recent advancements in single-token code completion, there is a niche of multi-token completion that is being actively studied in academia (Nijkamp et al., 2023b; Roziere et al., 2023; Fried et al., 2022; Li et al., 2023) and employed in products (Microsoft, 2024a; Amazon, 2024; TabNine, 2024; Codeium, 2024). Similar to single-token completion, multi-token completion can be obtained from a corresponding candidate provider. However, in contrast to standard code completion, it is usually shown as inline completion (also referred to as gray text completion, see Figure 1b) and thus usually provides only one suggestion.
@@
Most multi-token code completion solutions in both academia and industry tend to use language modelling approaches to generate suggestions. Specifically, they utilize Transformer-based neural networks (Vaswani et al., 2017) like GPT-2 (Radford et al., 2019) and LLaMA (Touvron et al., 2023). Such models usually have billions of parameters, implying gigabytes of memory footprint. Moreover, they are computationally intensive, so cloud-based multi-GPU servers are required to provide candidate suggestions.
@@
Figure 1. Code completion in the IDE: (a) standard lookup window, and (b) multi-token inline (gray text) completion.
2.1.2. Language modelling
@@
Over the past few years, Transformer-based neural networks started dominating in language modelling (Radford et al., 2019; Brown et al., 2020; Touvron et al., 2023). Specifically, models based solely on the decoder part of a Transformer (also known as autoregressive Transformers) are utilized for multi-token generation. A typical multi-token generation pipeline includes tokenization and sequence generation algorithms. Below, we briefly describe these parts for clearer understanding of the challenges we encountered when implementing FLCC.
@@
Tokenization. There exist many techniques for breaking down source sequences into tokens, including character-level, word-level, and subword-level tokenizations. The latter is typically used (Vaswani et al., 2017; Sennrich et al., 2015) with a Transformer-based model of any kind. Specifically, byte-pair encoding (BPE) (Radford et al., 2019; Touvron et al., 2023) is used for decoder-based autoregressive Transformers like GPTs and LLaMAs. This algorithm iteratively combines the most frequent pairs of adjacent tokens (bytes, characters, or character sequences) to form new, longer tokens. This process is repeated a predefined number of times or until a desired vocabulary size is reached. Then, during tokenization, the source sequence is transformed into a list of integers that correspond to an ordinal number of each token is the combined vocabulary.
@@
This brings up the problem of the search for the most probable token sequence. Since BPE vocabulary usually consists of tens of thousands of tokens, the generation of the distribution of the next step for every possible next token is not feasible. One of the popular solutions to this problem is the beam search algorithm (Huggingface, 2024b).
Figure 2. The general pipeline of a beam search algorithm.
@@
The overall scheme of the algorithm is presented in Figure 2. Beam search is a heuristic algorithm that focuses on traversing the graph of all possible sequences by expanding the node with the highest potential. This approach maintains only a fixed number of potential solutions at each step, called beams, or hypotheses. By focusing on the most promising hypotheses and discarding the less likely ones, beam search efficiently navigates through the graph. Usually, top $k$ most probable beams are supported at each step, and the next hypothesis score is obtained from the previous one, multiplying it by the probability of the next token. For Transformers, the procedure is implemented efficiently: the relations between context tokens can be computed once and reused afterward.
2.1.3. Release cycle of IDEs at JetBrains
@@
At JetBrains, an established IDE release cycle comprises 3 major releases per year, enumerated by year and year’s version in the following format: $YYYY.X$. Each release is preceded by a free early access program (EAP), allowing users to download experimental versions of the IDEs. Participation in the EAP includes user consent for the anonymized submission of usage logs to JetBrains. This data is crucial for our analysis, enabling us to determine the users’ experiences with new features. EAP releases are also used as a platform for A/B testing. Specifically, the EAP of the fall of 2023 was used to conduct the online evaluation for Full Line Code Completion.
2.2. Related Work
2.2.1. Datasets
@@
Many existing datasets for training code completion models, particularly for Python (Lu et al., 2021; Alon et al., 2020; 150k Python Dataset, 2024), have limitations due to licensing issues and quality concerns. Open-source code, often collected from platforms like GitHub, may have restrictive licenses or none at all (Golubev et al., 2020), which poses legal challenges (lic, 2022), as seen in GitHub Copilot’s (Microsoft, 2024a) usage. Furthermore, direct extraction of GitHub’s Python files yields many autogenerated and duplicated files (Kocetkov et al., 2022), leading to potential biases and data leakage issues in model training. A notable development to address these issues is the introduction of The Stack dataset (Huggingface, 2024d; Chen et al., 2021; Kocetkov et al., 2022) by BigCode (Huggingface, 2024a), a collaboration of over 350 researchers backed by HuggingFace and ServiceNow. This dataset, encompassing over 6TB of code in 358 languages, is extensively used for training large language models due to its permissive licensing and comprehensive coverage.
2.2.2. Code completion models
@@
In recent years, code completion became a major field of study for deep learning applications. In particular, many researchers and companies trained and released open-source models, resulting in the availability of training code, configuration specifics, and models’ weights.
@@
The used models are typically decoders of the Transformer architecture, including GPT and LLaMA architectures, which represent specific types of decoders. The most famous examples of such models are InCoder (Fried et al., 2022; et. al, 2024a), StarCoder (Li et al., 2023; BigCode, 2024a), CodeLLaMA (Roziere et al., 2023; Research, 2024), CodeGen and CodeGen-2 (Nijkamp et al., 2023b, a; SalesForce, 2024), CodeParrot (CodeParrot, 2024), and PolyCoder (Xu et al., 2022; et. al, 2024b). Despite the fact that the models are open-source, they are not suitable for local code completion engines like FLCC for several reasons. First of all, they are usually too big to ship to users and do not meet the memory footprint restrictions for desktop applications like IDEs. Secondly, the license for the model’s weights might not be permissive for commercial use. Last but not least, only specific architectures like GPT-2 and LLaMA have corresponding inference engines that allow to run them on an arbitrary end-user device fast enough.
2.2.3. Multi-token code completion products
@@
Lately, several products for multi-token completion appeared on the market, including Copilot (Microsoft, 2024a), TabNine (TabNine, 2024), Codeuim (Codeium, 2024), CodeWhisperer (Amazon, 2024), aiXcoder (aiXcoder, 2024), and others. Some of them remain free to use (e.g., Codeium and CodeWhisperer), while others require paid subscription. More importantly, all such products, except for TabNine and aiXcoder, offer only cloud-based solutions that might have certain restrictions for users behind the firewall, users with limited internet access, and users who have privacy concerns regarding sending their code to the cloud. Cloud-based solutions are unavailable not only for huge markets like China, but also for enterprise clients that want the source code to stay inside their network. Moreover, none of these solutions have a common API for inline completion, which makes them conflict with others and requires re-implementing the same functionality, which leads to poor UX and many confirmed bugs.
3. Approach
@@
  1. (1)
    The model and its execution engine must be local-based, i.e., operate fully on the user’s machine without sending anything over the internet. This is crucial because of privacy concerns and possible internet connection limitations.
  2. (2)
    Solution’s memory footprint must be feasible and take a reasonable part of the IDE memory footprint. This is necessary to ensure a convenient installation for the users and a feasible runtime, not overloading their machines for a single functionality. Insights from the IDE development teams give us a general upper bound of about 20-40% of the IDE memory footprint (might differ from product to product).
  3. (3)
@@
Figure 3. The overall pipeline of Full Line Code Completion.
@@
3.1. Data Collection
3.1.1. Dataset
We used the subset of The Stack (Huggingface, 2024d) — the StarCoder dataset (BigCode, 2024b), which we processed additionally to keep only permissive licenses. This completely fulfilled our needs in terms of handling licenses, as well as the filtering of duplicated files.
3.1.2. Train/test splitting
@@
After collecting the dataset, we found a representative set of repositories to be separated into the test part, so we can ensure that the model is not trained on them. This data is used later in a specialized evaluation setup described in Section 4.1. Additionally, to avoid potential data leaks, we ensure that all the forks of the given repository are placed in the same data segment. For the training of the Python model, the corresponding sizes of the train-validation-test sets of the data were 45.5 GB, 2.7 GB, and 5.4 GB.
3.2. Data Pre-processing

### v2

@@
Abstract
In recent years, several industrial solutions for the problem of multi-token code completion appeared, each making a great advance in the area but mostly focusing on cloud-based runtime and avoiding working on the end user’s device.
@@ Abstract
Our online evaluation shows that the usage of the tool leads to 1.3 times more Python code in the IDE being produced by code completion. The described solution was initially started with a help of researchers and was then bundled into all JetBrains IDEs where it is now used by millions of users. Thus, we believe that this work is useful for bridging academia and industry, providing researchers with the knowledge of what happens when complex research-based solutions are integrated into real products.
I Introduction
11footnotetext: These authors contributed equally to the work.
@@ I Introduction
Code completion is among the most attractive features in integrated development environments (IDEs) [1, 2], since it helps developers to avoid wasting time on manually typing all of the code characters and thus speeds up any software engineering activity. Depending on the specific IDE, the complexity of code completion (and thus its quality) varies a lot: it may just complete strings that a developer has already used in the current file, or it can use more complex approaches like static code analysis to predict possible tokens that a programmer may currently need. In any case, built-in completion engines usually only work in a single-token mode, i.e., only one identifier is generated (e.g., a variable or a method name).
@@
In the past years, several products for multi-token code completion have appeared on the market, including Copilot [3], TabNine [4], Codeium [5], Amazon CodeWhisperer [6], Sourcegraph Cody [7], aiXcoder [8], and others. Most of such tools are cloud-based, so they miss a wide niche of customers, which includes users beyond a firewall, as well as users with limited internet access. While TabNine and aiXcoder support fully local models and data querying, they lack integration features and do not implement code correctness checks. This still leaves an open market for code completion engines that are local-based and ensure a proper user experience in an IDE.
@@
In recent years, machine learning techniques were studied and applied for the different aspects of the completion problem [9, 10, 11, 12], including the generation of the completion sequences itself. Since source code can be interpreted as sequential data, natural language processing (NLP) methods have found their application for the task [13]. Specifically, Transformer-based [14] neural networks significantly outperform other architectures in the task of sequential data generation [15, 16, 17, 18].
@@
At JetBrains, we built the Full Line Code Completion (FLCC) feature [19] — a multi-token code completion engine that utilizes a Transformer-based neural network in its core and suggests only syntactically correct code. The feature is implemented in the form of a plugin for the IntelliJ Platform [20, 21] (an open-source platform for building IDEs), and is bundled into all JetBrains IDEs: IntelliJ IDEA [22], PyCharm Pro [23], WebStorm [24], and others, where it is enabled by default for various programming languages. The feature has been available for Python since version 2023.3 and then became available for other languages too in 2024.1 and 2024.2, now serving millions of users. It works fully locally, i.e. all computations, from code completion invocation to the showing of suggestion, happen on the end user’s device. Following modern industrial standards, the plugin is implemented in Kotlin, however, it utilizes an additional native server that is run locally and is implemented in C++. The whole model training pipeline is written mostly in Python, reusing the Kotlin pipeline only for the data processing step.
@@
In this paper, we describe how we built this Full Line Code Completion feature. Starting from the study of existing datasets, data collection, and the pre-processing with multiple tokenization specifics, we describe the details of the neural network training process and the end-user runtime, including a modified beam search algorithm, post-processing features, and user experience (UX) choices. While we cannot provide a full replication package of the described product feature, we share multiple conceptual insights we derived from our experience. The focus of the paper is on the Full Line Code Completion for Python, for which 100M parameters LLaMA-like [25] model was used, and inline completion (i.e., gray text completion) mode was implemented as a part of the IntelliJ Platform (see Figure 1b). Despite mainly focusing on Python, we share some evaluation results for other languages as well to provide a fuller picture. While preparing for the releases, we conducted rigorous offline evaluation, as well as online evaluation in the form of A/B testing. Our experiments showed that the ratio of completed code (i.e., the percentage of code symbols written with code completion among all written code) increased 1.3 times for Python users with Full Line Code Completion compared to users with standard code completion only.
@@
  • We provide an in-depth exploration of the implementation of Full Line Code Completion for 2024.1 & 2024.2 releases of IntelliJ-based IDEs.
@@
The rest of the paper is organized as follows. In Section II, we describe the related work in the field. Then, in Section III, we describe our approach in detail, highlighting various problems we encountered and how we solved them. Section IV presents the results of our A/B testing experiments, while Section V discusses these results and our lessons learned. Finally, in Section VI, we describe the threats to the validity of our work, and in Section VII, we conclude the paper.
II Background and Related Work
II-A Background
II-A1 Code completion
@@
Code completion, a prevalent feature across various IDEs [1, 2], is aimed at optimizing the work of a developer, not only by speeding up code typing but also by facilitating API exploration. Common in modern IDEs, single-token, or standard, code completion suggests individual tokens via a lookup window (see Figure 1a). This feature, relying on static code analysis, generates suggestions from multiple candidate providers, which include both built-in IDE modules and external plugins.
@@
In addition to recent advancements in single-token code completion, there is a niche of multi-token completion that is being actively studied in academia [26, 27, 28, 29] and employed in products [3, 6, 4, 5]. Similar to single-token completion, multi-token completion can be obtained from a corresponding candidate provider. However, in contrast to standard code completion, it is usually shown as inline completion (also referred to as gray text completion, see Figure 1b) and thus usually provides only one suggestion.
@@
Most multi-token code completion solutions in both academia and industry tend to use language modelling approaches to generate suggestions. Specifically, they utilize Transformer-based neural networks [14] like GPT-2 [15] and LLaMA [25]. Such models usually have billions of parameters, implying gigabytes of memory footprint. Moreover, they are computationally intensive, so cloud-based multi-GPU servers are required to provide candidate suggestions.
@@
Figure 1: Code completion in the IDE: (a) standard lookup window, and (b) multi-token inline (gray text) completion.
II-A2 Language modelling
@@
Over the past few years, Transformer-based neural networks started dominating in language modelling [15, 16, 25]. Specifically, models based solely on the decoder part of a Transformer (also known as autoregressive Transformers) are utilized for multi-token generation. A typical multi-token generation pipeline includes tokenization and sequence generation algorithms. Below, we briefly describe these parts for clearer understanding of the challenges we encountered when implementing FLCC.
@@
Tokenization. There exist many techniques for breaking down source sequences into tokens, including character-level, word-level, and subword-level tokenizations. The latter is typically used [14, 30] with a Transformer-based model of any kind. Specifically, byte-pair encoding (BPE) [15, 25] is used for decoder-based autoregressive Transformers like GPTs and LLaMAs. This algorithm iteratively combines the most frequent pairs of adjacent tokens (bytes, characters, or character sequences) to form new, longer tokens. This process is repeated a predefined number of times or until a desired vocabulary size is reached. Then, during tokenization, the source sequence is transformed into a list of integers that correspond to an ordinal number of each token is the combined vocabulary.
@@
This raises the problem of the search for the most probable token sequence. Since BPE vocabulary usually consists of thousands of tokens, the generation of the next step distribution for every possible next token is not feasible. One of the popular solutions to this problem is the beam search algorithm [31].
Figure 2: The general pipeline of a beam search algorithm.
@@
The overall scheme of the algorithm is presented in Figure 2. Beam search is a heuristic algorithm that focuses on traversing the graph of all possible sequences by expanding the node with the highest potential. This approach maintains only a fixed number of potential solutions at each step, called beams, or hypotheses. By focusing on the most promising hypotheses and discarding the less likely ones, beam search efficiently navigates through the graph. Usually, top $k$ most probable beams are supported at each step, and the next hypothesis score is obtained from the previous one, multiplying it by the probability of the next token. For Transformers, this procedure has an efficient implementation: the relations between context tokens can be computed once and reused afterward.
II-A3 Release cycle of IDEs at JetBrains
@@
At JetBrains, an established IDE release cycle comprises 3 major releases per year, enumerated by year and year’s version in the following format: $YYYY.X$. Each release is preceded by a free early access program (EAP), allowing users to download experimental versions of the IDEs. Participation in the EAP includes user consent for the anonymized submission of usage logs to JetBrains. This data is crucial for our analysis, enabling us to determine the users’ experiences with new features. EAP releases are also used as a platform for A/B testing. The EAPs of versions 2023.3, 2024.1, and 2024.2 were used to conduct the online evaluation for Full Line Code Completion.
II-B Related Work
II-B1 Datasets
@@
Many existing datasets for training code completion models, particularly for Python [32, 33, 34], have limitations due to licensing issues and quality concerns. Open-source code, often collected from platforms like GitHub, may have restrictive licenses or none at all [35], which poses legal challenges [36], as seen in GitHub Copilot’s [3] usage. Furthermore, direct extraction of GitHub’s Python files yields many autogenerated and duplicated files [37], leading to potential biases and data leakage issues in model training. A notable development to address these issues is the introduction of The Stack dataset [38, 9, 37] by BigCode [39], a collaboration of over 350 researchers backed by HuggingFace and ServiceNow. This dataset, encompassing over 6TB of code in 358 languages, is extensively used for training large language models due to its permissive licensing and wide coverage.
Figure 3: The overall pipeline of Full Line Code Completion.
II-B2 Code completion models
@@
In recent years, code completion became a major field of study for deep learning applications. Many researchers and companies trained and released open-source models, resulting in the availability of training code, configuration specifics, and models’ weights.
@@
The employed models are typically decoders of the Transformer architecture, including GPT and LLaMA architectures representing specific types of decoders. The most famous examples of such models are InCoder [28, 40], StarCoder [29, 41], CodeLLaMA [27, 42], CodeGen and CodeGen-2 [26, 43, 44], CodeParrot [45], and PolyCoder [46, 47]. Despite the fact that the models are open-source, they are not suitable for local code completion engines like FLCC for several reasons. First of all, they are usually too big to ship to users and do not meet the memory footprint restrictions for desktop applications like IDEs. Secondly, the license for the model’s weights might not be permissive for commercial use. Last but not least, only specific architectures like GPT-2 and LLaMA have corresponding inference engines that allow to run them on an arbitrary end-user device fast enough.
II-B3 Multi-token code completion products
@@
Lately, several products for multi-token completion appeared on the market, including Copilot [3], TabNine [4], Codeuim [5], CodeWhisperer [6], aiXcoder [8], Sourcegraph Cody [7], and others. Some of them remain free to use (e.g., Codeium and CodeWhisperer), while others require paid subscription. More importantly, all such products, except for TabNine and aiXcoder, offer only cloud-based solutions that might have certain restrictions for users behind the firewall, users with limited internet access, and users who have privacy concerns regarding sending their code to the cloud. Cloud-based solutions are unavailable not only for huge markets like China, but also for enterprise clients that want the source code to stay inside their network. Moreover, none of these solutions have a common API for inline completion, which makes them conflict with others and requires re-implementing the same functionality, which leads to poor UX and many confirmed bugs.
III Approach
@@
  1. 1.
    The model and its execution engine must be local-based, i.e., operate fully on the user’s machine without sending anything over the internet. This is crucial because of privacy concerns and possible internet connection issues.
  2. 2.
    Solution’s memory footprint must be feasible and take a reasonable part of the IDE memory footprint. This ensures a convenient installation and a feasible runtime, not overloading the users’ machines for a single functionality. Insights from the IDE development teams give us a general upper bound of about 20–40% of the IDE memory footprint (might differ between products).
  3. 3.
@@
III-A Data Collection
III-A1 Dataset
For Python, we used the subset of The Stack [38] — the StarCoder dataset [48], which we additionally processed to keep only permissive licenses. This completely fulfilled our needs in terms of handling licenses, as well as the filtering of duplicated files.
III-A2 Train/test splitting
@@
After collecting the dataset, we found a representative set of repositories to be separated into the test part, thus ensuring that the model is not trained on them. This data is used later in a specialized offline evaluation setup described in Section IV-A. Additionally, to avoid potential data leaks, we ensure that all the forks of the given repository are placed in the same data segment. For the training of the Python model, the corresponding proportions of the train-validation-test sets of the data were 80%-5%-15%, adding up to slightly more than 50 GB of data.
III-B Data Pre-processing